# Space-Time Graphs of Convex Sets for Multi-Robot Motion Planning

### v1

@@ 8             $\mathcal{X}_{v_{i}}\leftarrow\mathcal{X}_{v_{i}}\setminus\text{outside}(L_{i},s)$
Vertex-States Sequence Construction: Before applying ECD to reserve a piecewise linear trajectory $\tau$ on a given graph $\mathcal{G}$, we must construct a sequence of vertex-states tuples $\{(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{k}$ such that each segment $\mathbf{x}_{i}\rightarrow\mathbf{y}_{i}$ lies entirely within one single $\mathcal{X}_{v_{i}}$. This construction is particularly relevant in our prioritized planning frameworks (Sec. IV-B) and for dynamic obstacles, where ECD reserves trajectories for collision avoidance. Three cases arise: (1) If $\tau$ is reconstructed from an ST-GCS solution $(\Phi,\mathbf{x},\mathbf{y})$ computed on the same graph $\mathcal{G}$, each segment $\mathbf{x}\rightarrow\mathbf{y}$ naturally maps to a unique convex set of $\mathcal{G}$. (2) If a segment $\mathbf{x}\rightarrow\mathbf{y}$ intersects more than one convex set of $\mathcal{G}$, we subdivide it at each boundary so that each sub-segment lies in exactly one convex set. (3) If a convex set $\mathcal{X}_{v}$ contains multiple segments $\mathbf{x}\rightarrow\mathbf{y}$ (non-overlapping in time), we slice $\mathcal{X}_{v}$ by the planes $t=\mathbf{x}.t$ and $t=\mathbf{y}.t$, ensuring each resulting subset contains at most one segment. Note that, when our ST-GCS is used with our prioritized planning frameworks, only cases (1) and (2) occur.
@@ V Experiments
This section presents our experimental results on an Apple® M4 CPU machine with 16GB RAM. We evaluate RP+ST-GCS and PBS+ST-GCS against three baseline methods in 2D mobile robot domains. All methods are implemented in Python, and our ST-GCS MICP is solved using the Drake [32] library with the Mosek [^4] https://www.mosek.com/ solver. All the source code and numerical results are available on GitHub [^5] https://github.com/reso1/stgcs.
@@ V-A Experiment Setup
Parameters: For empty, we set a maximum velocity limit of $0.5$ for each space dimension; For empty and complex, we set a maximum velocity limit of $1.0$ for each spatial dimension. We set a runtime limit of 150 seconds for all methods. SP+T-PRM allocates $150/n$ seconds for each robot’s PRM construction (omitting its shortest-path finding runtime), and SP+ST-RRT allocates $150/n$ seconds for each single-robot ST-RRT planning. For ST-GCS MICP solving on any graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ using the heuristic approach (see Sec. IV-B), we limit the number of graph paths sampled to $1e^{3}\times\log|\mathcal{E}|$.
@@ V-B Results and Analysis
Success Rates and Runtimes: PBS+ST-GCS solves all instances on empty and complex, with average runtimes consistently under 1 second and 10 seconds on average, respectively—often orders of magnitude faster than sampling-based methods. Adding dynamic obstacles on simple significantly enlarges the initial space-time graph (via ECD), which can slow ST-GCS solving; nevertheless, PBS+ST-GCS still achieves the highest success rates overall. RP+ST-GCS randomly explores different priority orders until timeout generally attains higher success rates than SP+ST-GCS which uses only a fixed priority order. Among sampling-based methods, SP+ST-RRT outperforms SP+T-PRM on empty and simple, likely due to the inherent faster tree-based exploration rooted at the starts and goals. However, on complex with many spatial corridors, the more global random exploration of T-PRM yields higher success rates than ST-RRT. Still, both fail when $n>2$ on complex. Comparing all SP-based methods, SP+ST-GCS typically runs faster than the two sampling-based methods and achieves higher success rates in settings with more spatial corridors. However, its success rates are lower for larger $n$, due to the heuristic MICP solver failing more frequently, as analyzed in the ablation study below.
@@
Ablation—Effectiveness of PBS for ST-GCS: Our raw data indicates that large space-time graphs (over 1000 edges) resulting from numerous dynamic obstacles can impede the heuristic MICP solver under limited path-sampling budgets. Fig. 5 demonstrates that PBS mitigates this issue by reserving only conflicting trajectories deemed high-priority, controlling graph growth more effectively than SP+ST-GCS and RP+ST-GCS (which both rely on total priority orders with sequential ECD, thus expanding the graph faster as $n$ grows).
@@ V-C Case Study
We highlight two examples of ST-GCS and PBS+ST-GCS. Figure 7 demonstrates ST-GCS finding a safe shortcut through dynamic obstacles, where ST-RRT struggles to connect sample points to exploring trees, and T-PRM fails to find a feasible solution even after 20k samples. Figure 8 compares PBS+ST-GCS against SP+ST-RRT for 8 robots on simple. PBS+ST-GCS lets them move through the center congested region simultaneously with minimal waiting or detours, whereas SP+ST-RRT schedules them one by one, creating more zig-zagging and delays.
Figure 7: Single-robot motion planning on simple. A robot starts from the bottom-left at $t=0$ and must reach the top-right while avoiding collisions with four dynamic obstacles (black disks). ST-RRT solution reaches goal at $t=8.29$ with $30$-second runtime, while our ST-GCS solution reaches goal at $t=5.90$ with $0.6$-second runtime.
Figure 8: MRMP on simple where 8 robots must switch their positions. Start and goal positions of robot $i$ are denoted as $s_{i}$ and $g_{i}$, respectively. (a)(b) SP+ST-RRT and PBS+ST-GCS solutions visualized in space-time. (c)(d) Magnified views of the center congested region at time $t=2.5$.
@@ VI Conclusions & Future Work
We presented ST-GCS, a time-optimal deterministic approach to MRMP that addresses the inherent limitations of sampling-based methods in spatiotemporal settings. By extending the GCS formulation to a space-time domain, ST-GCS systematically covers spatiotemporal bottlenecks using collision-free convex sets. Our ECD algorithm further enables straightforward reservation of piecewise linear trajectories of dynamic obstacles for collision avoidance. We integrated ST-GCS and ECD into prioritized planning frameworks and demonstrated through extensive experiments that our approach consistently outperforms state-of-the-art sampling-based methods in both success rates and solution quality, often with orders-of-magnitude faster runtimes, especially in challenging scenarios such as narrow corridors and crowded environments. Future work includes investigating robust pruning strategies for managing the ST-GCS graph size when numerous trajectories are reserved via ECD, generalizing ECD to higher-dimensional spaces and more complex obstacle models, and extending ST-GCS with richer motion models with nonholonomic and dynamic constraints.

### v2

@@ 8             $\mathcal{X}_{v_{i}}\leftarrow\mathcal{X}_{v_{i}}\setminus\text{outside}(L_{i},s)$
Vertex-Segment Sequence Construction: We describe how to construct a sequence $S=\{(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{k}$ of vertex-segment tuples from a piecewise linear trajectory $\tau$ on $\mathcal{G}$. This construction is particularly relevant in our prioritized planning frameworks (Sec. IV-B) and for dynamic obstacles, where ECD reserves trajectories for collision avoidance. Three cases arise: (1) If $\tau$ is reconstructed from an ST-GCS solution computed on the same graph $\mathcal{G}$, each segment of $\tau$ naturally maps to a unique convex set $\mathcal{X}_{v}$ of $\mathcal{G}$, and potentially to the adjacent convex sets of $\mathcal{X}_{v}$ if the segment lies in the boundaries of $\mathcal{X}_{v}$. (2) If a segment of $\tau$ intersects more than one convex set of $\mathcal{G}$, we subdivide it at each boundary so that each sub-segment lies in exactly one convex set. (3) If a convex set $\mathcal{X}_{v}$ contains multiple segments $\mathbf{x}_{j}\rightarrow\mathbf{y}_{j}$ (non-overlapping in time), we slice $\mathcal{X}_{v}$ by the planes $t=\mathbf{x}_{j}.t$ and $t=\mathbf{y}_{j}.t$, ensuring each resulting subset contains at most one segment. We adopt a unified procedure to construct the sequence $S$, considering the above three cases. For each linear segment in $\tau$, we check whether it intersects with each convex set of $\mathcal{G}$444 Our implementation leverages bounding boxes for the convex sets to do coarse collision checking before exact collision checking for efficiency. . If intersecting, we collect the vertex $v_{i}$ and the two endpoints $\mathbf{x}_{i},\mathbf{y}_{i}$ of the intersecting sub-segment into sequence $S$.
@@ V Experiments
This section presents our experimental results on an Apple® M4 CPU machine with 16GB RAM. We evaluate RP+ST-GCS and PBS+ST-GCS against three baseline methods in 2D mobile robot domains. All methods are implemented in Python, and our ST-GCS program is solved using the Drake [34] library with the Mosek [^5] https://www.mosek.com/ solver. The source code and numerical results are publicly available on https://github.com/reso1/stgcs. More detailed visualizations and simulation videos of our approach can be found at https://sites.google.com/view/stgcs.
@@ V-A Experiment Setup
Parameters: For empty, we set a maximum velocity limit of $0.5$ for each space dimension; For empty and complex, we set a maximum velocity limit of $1.0$ for each spatial dimension. We set a runtime limit of 150 seconds for all methods. SP+T-PRM allocates $150/n$ seconds for each robot’s PRM construction (omitting its shortest-path finding runtime), and SP+ST-RRT allocates $150/n$ seconds for each single-robot ST-RRT planning. For ST-GCS solving on any graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ using the heuristic approach (see Sec. IV-B), we limit the number of graph paths sampled to $1e^{3}\times\log|\mathcal{E}|$.
@@ V-B Results and Analysis
Success Rates and Runtimes: PBS+ST-GCS solves all instances on empty and complex, with average runtimes consistently under 1 second and 10 seconds on average, respectively—often orders of magnitude faster than sampling-based methods. Adding dynamic obstacles on simple significantly enlarges the initial space-time graph (via ECD), which can slow ST-GCS solving; nevertheless, PBS+ST-GCS still achieves the highest success rates overall. RP+ST-GCS randomly explores different priority orders until timeout generally attains higher success rates than SP+ST-GCS which uses only a fixed priority order. Among sampling-based methods, SP+ST-RRT outperforms SP+T-PRM on empty and simple, likely due to the inherent faster tree-based exploration rooted at the starts and goals. However, on complex with many spatial corridors, the more global random exploration of T-PRM yields higher success rates than ST-RRT. Still, both fail when $n>2$ on complex. Comparing all SP-based methods, SP+ST-GCS typically runs faster than the two sampling-based methods and achieves higher success rates in settings with more spatial corridors. However, its success rates are lower for larger $n$, due to the heuristic solver failing more frequently, as analyzed in the ablation study below.
@@
Ablation—Effectiveness of PBS for ST-GCS: Our raw data indicates that large space-time graphs (over 1000 edges) resulting from numerous dynamic obstacles can impede the heuristic solver under limited path-sampling budgets. Fig. 5 demonstrates that PBS mitigates this issue by reserving only conflicting trajectories deemed high-priority, controlling the graph size growth more effectively than SP+ST-GCS and RP+ST-GCS (which both rely on total priority orders with sequential ECD, thus expanding the graph faster as $n$ grows).
@@ V-C Case Study
We highlight two examples demonstrating the advantages of PBS+ST-GCS over SP+ST-RRT. Fig. 7 demonstrates that PBS+ST-GCS identifies a safe shortcut through dynamic obstacles, producing a time-optimal MRMP solution. Fig. 8 shows that PBS+ST-GCS enables robots to traverse the congested central region simultaneously with minimal waiting or detours, while SP+ST-RRT generates zigzagging trajectories and significant delays (e.g., robot 5 is delayed until other robots nearly reach their goals).
Figure 7: Two robots exchange positions with four dynamic obstacles (black).
Figure 8: Eight robots exchange positions with one another on simple.
@@ VI Conclusions & Future Work
We presented ST-GCS, a time-optimal deterministic approach that addresses the inherent limitations of sampling-based methods in spatiotemporal settings. By extending the GCS formulation to a space-time domain, ST-GCS systematically covers spatiotemporal bottlenecks using collision-free convex sets. Our ECD algorithm further enables straightforward reservation of piecewise linear trajectories of dynamic obstacles for collision avoidance. We integrated ST-GCS and ECD into prioritized planning frameworks for MRMP. We demonstrated through extensive experiments that our approach consistently outperforms state-of-the-art sampling-based methods in both success rates and solution quality, often with orders-of-magnitude faster runtimes, especially in challenging scenarios such as narrow corridors and crowded environments. Future work includes investigating robust pruning strategies for managing the ST-GCS graph size when numerous trajectories are reserved via ECD, generalizing to high-dimensional nonlinear configuration spaces and more complex obstacle models, and extending ST-GCS with richer motion models with nonholonomic and dynamic constraints.
Acknowledgement
This work was supported by the NSERC under grant number RGPIN2020-06540 and a CFI JELF award. We thank the anonymous reviewers for their constructive feedback that helped us improve this paper.